Gamma rays make sterilization easier they are used in autoclaves. They emits electromagnetic wave of wavelength 0.6*10^-9m determine the period of wave lenghth
Deduce the frequency of this wave
1
Expert's answer
2020-03-16T12:59:59-0400

The period of wave and its frequency related to the wavelength via the propagation speed. For gamma rays wich are the  kind of electromagnetic wave with speed of propagation "c=3\cdot 10^8 ms^{-1}" .

For frequency of gamma rays we have "f=\frac{c}{\lambda}=\frac{3\cdot 10^8 ms^{-1}}{0.6\cdot 10^{-9}m}=5\cdot10^{17}Hz" .


For period of gamma rays we have "T=\frac{\lambda}{c}=\frac{1}{f}=2\cdot 10^{-18}s" .

Answer: Frequency of gamma rays with wavelength "\lambda =0.6\cdot 10^{-9}m" is "f=5\cdot10^{17}Hz" and period is "T=2\cdot 10^{-18}s".